# Q*: Improving Multi-step Reasoning for LLMs with Deliberative Planning

### v1

@@ 5.1 Experimental Settings
Implementation Details. The implementation of Q* method mainly include three steps: 1) Q-value estimation. As discussed in Section 4.1, we propose several ways for estimating Q values and will illustrate the implementation details in the next paragraph; 2) Aggregated utility for each step (cf. Eq. (4)). For GSM8K dataset, we adopt a process reward model $\mathcal{R}_{P}$ trained on PRM800K [22] to provide intermediate signal for each reasoning step, and use $\min$ as the aggregation function; For MATH dataset, we set $g(s)=0$ for all state $s$ for fairness because PRM800K contains samples constructed from MATH testing set and the trained PRM could potentially have seen the correct answer of testing problems; For MBPP dataset, we tokenize the code generated so far with function tokenize.generate_tokens and give a penalty of -0.5 if TokenError is raised, which is often the case that there are mismatched delimiters (e.g., parentheses, quotation marks) and invalid indention in the code. We use $[-1]$ as the aggregation function to cancel the previous penalties since the code is generated on-the-fly and mismatched delimiters may be fixed in subsequent steps. 3) A* planning. For GSM8K and MATH datasets, we treat a single line outputted by the LLM as an action, while the action in MBPP is defined as a code snippet with 24 tokens when planning. Besides, we set $\lambda=1$ when computing $f$-values and expand a state with $K=6$ actions. Finally, following the common practise of Best-of-$N$, in all experiments we perform planning to collect $N=6$ trajectories for each question, and select the one with the maximum $f$-value as the final result for evaluation.
[…]
@@ 5.2 Experimental Results
GSM8K. For the comparison on GSM8K dataset, we select Llama-2-7b [44] as our base model, whose accuracy can achieve 65.2% after finetuning on MetaMath [5]. Then, we treat Llama-2-7b finetuned on MetaMath as policy $\pi_{\theta}$, and perform rollout to collect Q-value labels for training Q-value model (QVM). For aggregated utility, we train a process reward model (PRM) on PRM800K [22] to provide intermediate signal for each reasoning step. With a PRM/QVM in hand, traditional methods tend to treat either of them as a verifier the select the Best-of-$N$ trajectory or utilize them to perform PPO training of RLHF. As the results shown in Table 2, we can find that with the same PRM/QVM, using it for verification performs significantly better than using it for alignment. Further, in the comparison of planning-based methods, we can find that with the same QVM, Q* method with constant aggregated utility can still outperform Best-of-$N$ method. With the PRM trained on PRM800K determining whether the intermediate reasoning steps are correct, Q* method that combines PRM and QVM achieves the best performance among the methods based on the Llama-2-7b model, surpassing the performance of close-sourced ChatGPT-turbo [45].
MATH. As the results shown in Table 3, considering the weak performance of Llama-2-7b fine-tuned on MetaMath for the MATH dataset, we seek for two other stronger base models to evaluate the effectiveness of the Q* method. The one is Llama-2-7b fine-tuned on Synthetic Data [46], which is constructed following the instruction of scaling up the SFT data, and achieves 41.9% accuracy on MATH dataset, approaching the performance of GPT-4 [47]. Another base model is DeepSeek-Math-7b [48], which could be the most powerful open-source 7b model for math reasoning, achieving 50.8% accuracy in our evaluation. From the results shown in the second and third blocks of Table 3, we found that the Q* method can lead to further performance improvements compared to the Best-of-$N$ method. Additionally, the performance of Q* based on DeepSeek-Math-7b has already surpassed a series of closed-source models on the leaderboard of MATH dataset  [^1] https://paperswithcode.com/sota/math-word-problem-solving-on-math, such as Gemini Ultra (4-shot).
@@
MBPP. As for the comparison on MBPP dataset, we also choose the most powerful open-source LLM in the aspect of code generation, specifically CodeQwen1.5-7b-Chat, as our base model for evaluating the effectiveness of Q*. Following a similar procedure of math reasoning, we train a QVM for Q-value estimation and manually construct heuristic function as described in the previous part of implementation details. From the results shown in Table 4, we can find that Q* can still outperform Best-of-$N$ method in the aspect of code generation, and achieve 77.0% in MBPP dataset, which is still a promising performance in the leaderboard of MPBB  [^2] https://paperswithcode.com/sota/code-generation-on-mbpp.
@@ 6 Conclusion
Solving challenging multi-step reasoning problems requires LLMs to perform in-depth deliberation beyond auto-regressive token generation. In this paper, we present Q*, a general, versatile and agile deliberation framework for LLMs. Unlike existing deliberation methods which need extensive expertise to design a utility function for each specific task, Q* relies on ground-truth solely to train value model and can be easily applied to various reasoning tasks without modification. Moreover, by leveraging plug-and-play Q-value models as heuristic function, Q* can effectively guide LLMs to solve various tasks without fine-tuning LLMs beforehand, which avoid potential performance degeneration on other tasks. Finally, our Q* is agile because we consider only a single step each time rather than complete rollouts (e.g., simulation in MCTS). Extensive empirical evaluations on math reasoning and code generation tasks confirm the superiority of our method.

### v2

@@ 5.1 Experimental Settings
Implementation Details. The implementation of Q* method mainly includes three steps: 1) Q-value estimation. As discussed in Section 4.1, we propose several ways for estimating Q values and will illustrate the implementation details in the next paragraph; 2) Utility aggregation for each step (cf. Eq. (5)). For GSM8K dataset, we adopt a process reward model (PRM) trained on PRM800K [22] to model $\mathcal{R}_{P}$ to provide an intermediate signal for each reasoning step, and use $\min$ as the aggregation function; For MATH dataset, we set $g(s_{t})=0$ for all passed states $\{s_{i}\}_{i=1}^{t}$ in each trajectory for fairness, because PRM800K contains data samples constructed from MATH testing set and there is a potential risk of data leakage; For MBPP dataset, we tokenize the code generated so far with function tokenize.generate_tokens and give a penalty of -0.5 if TokenError is raised, which is often the case that there are mismatched delimiters (e.g., parentheses, quotation marks) and invalid indention in the code. We use $[-1]$ as the aggregation function to cancel the previous penalties since the code is generated on-the-fly and mismatched delimiters may be fixed in subsequent steps. 3) A* planning. For GSM8K and MATH datasets, we treat a single line outputted by the LLM as an action, while the action in MBPP is defined as a code snippet with 24 tokens when planning. Besides, in all experiments, we set $\lambda=1$ when computing $f$-values and expand a state with $K=6$ actions at each reasoning step. Finally, following the common practice of Best-of-$N$, we perform planning to collect $N=6$ trajectories for each question, and select the one with the maximum $f$-value as the final result for evaluation.
For Q-value estimation, in our practice, we find that learning from rollout could be the most effective and robust way to collect precise Q-value labels. Specifically, given the prompt of questions, we will firstly perform random rollout to obtain complete trajectories with LLM, denoted as $\pi_{\theta}$, under the setting of high temperature, e.g., $\tau=0.9$ for math reasoning and $\tau=0.2$ for code generation, and split each trajectory into a series of states according to the $\n\𝑛\textbackslash n$”. Then, for each state-action pair in a trajectory, denoted as $(s_{t},a_{t})$, we can perform random rollout or MCTS with the same LLM to collect a pool $P$ of trajectories, and then select the best reasoning path with the highest accumulated rewards to construct the corresponding Q-value label of the current state-action pair. Note that the reward $\mathcal{R}(s_{t},a_{t})$ is given as 1 only if the obtained math numerical answer is correct or the program passes all test cases, indicating that the Q value of a state-action pair can be 1 only if it has the potential to generate a trajectory containing the correct answer.
@@ 5.2 Experimental Results
GSM8K. For the comparison on GSM8K dataset, we select Llama-2-7b [45] as our base model, whose accuracy can achieve 65.2% after finetuning on MetaMath [5]. Then, we treat Llama-2-7b finetuned on MetaMath as policy $\pi_{\theta}$, and perform random rollout to collect Q-value labels for training Q-value model (QVM). For utility aggregation, we train a process reward model (PRM) on PRM800K [22] to provide intermediate signal for each reasoning step. With PRM and QVM in hand, traditional methods tend to treat either of them as a verifier to select the Best-of-$N$ trajectory or utilize them to perform PPO training of RLHF. As the results shown in Table 2, we can find that with the same PRM/QVM, using it for verification performs significantly better than using it for alignment. Further, in the comparison of planning-based methods, we can find that with the same QVM, Q* method with constant aggregated utility can still outperform Best-of-$N$ method. With the PRM trained on PRM800K determining whether the intermediate reasoning steps are correct, Q* method that combines PRM and QVM achieves the best performance among all methods based on the same LLM, helping Llama-2-7b surpass the performance of close-sourced ChatGPT-turbo [46] and reaching an accuracy of 80.8%.
MATH. As the results shown in Table 3, considering the weak performance of Llama-2-7b fine-tuned with MetaMath for the MATH dataset, we seek for two other stronger LLMs to evaluate the effectiveness of our Q* method. One is Llama-2-7b fine-tuned on Synthetic Data [47], which is constructed following the instruction of scaling up the SFT data, and achieves 41.9% accuracy on MATH dataset, approaching the performance of GPT-4 [48]. The other base model is DeepSeek-Math-7b [49], which could be the most powerful open-source 7b model for math reasoning on MATH dataset, achieving 50.8% accuracy in our evaluation. From the results shown in the second and third blocks of Table 3, we can find that Q* can still lead to further performance improvement compared to the Best-of-$N$ method on either of base models. Additionally, it is noteworthy that the performance of DeepSeek-Math-7b enhanced with Q* has already surpassed a series of closed-source models on the leaderboard of MATH dataset  [^1] https://paperswithcode.com/sota/math-word-problem-solving-on-math, such as Gemini Ultra (4-shot), reaching an accuracy of 55.4% .
@@
MBPP. As for the comparison on MBPP dataset, we also choose the most powerful open-source LLM in the aspect of code generation, specifically CodeQwen1.5-7b-Chat, as our base model for evaluating the effectiveness of Q*. Following a similar procedure of math reasoning, we train a QVM for Q-value estimation and manually construct the utility function as described in the previous part of implementation details (tokenize the code generated so far with function tokenize.generate_tokens and give a penalty of -0.5 if TokenError is raised). From the results shown in Table 4, we can find that Q* can still outperform Best-of-$N$ method in the aspect of code generation, and help CodeQwen1.5-7b-Chat to achieve 77.0% accuracy on MBPP dataset, which is also a promising performance in the leaderboard of MPBB  [^2] https://paperswithcode.com/sota/code-generation-on-mbpp.
@@ 6 Conclusion
Solving challenging multi-step reasoning problems requires LLMs to perform in-depth deliberation beyond auto-regressive token generation. In this paper, we present Q*, a general, versatile and agile deliberation framework for LLMs. Unlike existing deliberation methods which need extensive expertise to design a utility function for each specific task, Q* relies on ground-truth solely to train value model and can be easily applied to various reasoning tasks without modification. Moreover, by leveraging plug-and-play Q-value models as the heuristic function, Q* can effectively guide LLMs to solve various tasks without fine-tuning LLMs beforehand, which avoids potential performance degeneration on other tasks. Finally, our Q* is agile because we consider only a single step each time rather than complete rollouts (e.g., simulation in MCTS). Extensive empirical evaluations on math reasoning and code generation tasks confirm the superiority of our method.